# StylizedGS: Controllable Stylization for 3D Gaussian Splatting

### v1

@@ 5 Discussion and Conclusion
Limitations. Although our method achieves efficient and controllable 3DGS stylization, it still has some limitations. First, the geometric artifacts from the original 3DGS reconstruction may impact the quality of the final stylized scenes. Although our 3DGS filter can eliminate some floaters, it cannot eliminate all geometric artifacts. This can rely on some improved 3DGS reconstruction methods that add geometric constraints, such as SuGaR [53]. Another point is that our optimization-based method cannot achieve instant stylization. However, this does not compromise our method’s stylization effect and practical application as the optimization time is controlled at about 1 minute.
In this paper, we propose StylizedGS, the first controllable 3D scene stylization method based on 3D Gaussian Splatting representation. After the 3DGS reconstruction, the proposed GS filter minimizes the impact of floaters in the reconstruction, which is crucial for the final stylization effect. Then we propose to adopt the nearest-neighbor feature matching style loss to optimize both geometry and color parameters of 3D Gaussians to capture the detailed style feature and facilitate the stylization of the 3D scene. We further introduce a depth preservation loss for regularization to preserve the overall structure during the stylization. Moreover, we design controllable ways for three perceptual factors: color, stylization scale, and spatial regions, so our method supports specific and diverse user control. Qualitative and quantitative experiments demonstrate that our method outperforms existing 3D stylization methods in terms of effectiveness and efficiency.

### v2

@@ 5 Discussion and Conclusion
Figure 15: Limitation. The geometric artifacts from the original 3DGS reconstruction may impact the quality of the final stylized scenes.
Limitations. Although our method achieves efficient and controllable 3DGS stylization, it still has some limitations. First, the geometric artifacts from the original 3DGS reconstruction may impact the quality of the final stylized scenes, as shown in Fig. 15. Although our filter-based refinement can eliminate some floaters, it cannot eliminate all geometric artifacts. In the future, we will incorporate some improved 3DGS reconstruction methods that add geometric constraints, such as SuGaR [71], to address this issue. In addition, our optimization-based method cannot achieve instant stylization, However, this does not compromise the effectiveness of our method or its practical value, as the optimization process typically takes only about 1 minute.
In this paper, we introduce StylizedGS, the first controllable 3D scene stylization method based on the 3D Gaussian Splatting representation. Following 3DGS reconstruction, our proposed filter-based refinement minimizes the impact of floaters in the reconstruction, which is crucial for achieving the desired stylization effect. Additionally, we propose the adoption of nearest-neighbor feature matching style loss to optimize both geometry and color parameters of 3D Gaussians, enabling the capture of detailed style features and facilitating 3D scene stylization. We further introduce a depth preservation loss for regularization to maintain the overall structure during stylization. Moreover, we design controllable methods for three perceptual factors: color, stylization scale, and spatial regions, providing users with specific and diverse control options. Qualitative and quantitative experiments demonstrate that our method outperforms existing 3D stylization methods in terms of effectiveness and efficiency.